at = 0.02 m/s^2 where t is the time in seconds measured from rest. Imagine starting @ C. Calculate the velocity and aceleration at point B.
120 250 m
Answer #1

oiven, New dt dt ナc. at, tzo, 0 ナゾ 0.01 x t乙 nlonw, d t at at t20, 20 3

Now,

Distance to be traveled on straight track, d = 250 m

Now,

radius of the curvature, r = 50 m

From figure we can see angle covered is 120^{circ}

Thus, distance to be traveled on curved path, d' = (120/360)*2*pi*r = 1/3*2*3.14*50

                                                                              = 104.66 m

Thus, total distance traveled, x = d + d' = 250 + 104.66

                                                   = 354.66 m

As we have,

=> x = 0.0033 * t3

=> 354.66 = 0.0033 * t3

=> t3 = 354.66 / 0.0033 = 107472.7

=> t = V107472.7 = 47.54 s

Now,

v = 0.01 * t2 = 0.01 * 47.542

   = 22.6 m/s

and

a = 0.02*t = 0.02 * 47.54

   = 0.95 m/s2